OpenStudy (anonymous):

HELPP Felipe's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Felipe $5.65 per pound, and type B coffee costs $4.55 per pound. This month, Felipe made 148 pounds of the blend, for a total cost of $761.40 . How many pounds of type B coffee did he use?

OpenStudy (anonymous):

two variables involved....one equation thus cannot be solved by normal algebra and try to do that by hit and trial

OpenStudy (anonymous):

How many pounds of type B?

OpenStudy (skullpatrol):

Felipe made pounds of the blend? How many?

OpenStudy (anonymous):

148

OpenStudy (anonymous):

no its there

OpenStudy (skullpatrol):

Yeah, NOW

OpenStudy (anonymous):

ohh darn

OpenStudy (anonymous):

well

OpenStudy (anonymous):

let the pounds of A be x and for B be y 5.65x + 4.55y= 761.40 x+y= 148

OpenStudy (skullpatrol):

($5.65 per pound)*A + ($4.55 per pound)*B = for a total cost of $761.40 A + B = 148 pounds of the blend

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

74?

OpenStudy (skullpatrol):

If you solve for A too, you can check your own answer, right?

OpenStudy (anonymous):

so is it 74

OpenStudy (skullpatrol):

@kaylee.carmestro are you not interested enough to solve for B and check your own answer?

OpenStudy (anonymous):

yesss

OpenStudy (skullpatrol):

Check your results with the words of the problem, Give the answer. You will gain confidence that way too :-)